OpenStudy (anonymous):

What is the nth term of this sequence? 1/3, 1/2, 3/5, 2/3, 5/7 A.) n / n-2(n+3) B.) n / n+2 C.) n / 2n+1 D.) n+2 / N-1 Please help! Thank you<3 xoxo

OpenStudy (anonymous):

Heeelp ;~; Pweeeeeezzzz.

OpenStudy (anonymous):

you can just plug in the sequence indices to each choice until you get one that matches. \(a_1=\dfrac{1}{3}\\a_2=\dfrac{1}{2}\\a_3=\dfrac{3}{5}\\....\)

OpenStudy (anonymous):

Let me try..

OpenStudy (trojanpoem):

Use the arithmetic rule

OpenStudy (trojanpoem):

r= 1/2 - 1/3 = 3/6 - 2/6 = 1/6 a (first in the sequence) = 1/3 nth term = a + r( n - 1) = 1/3 + 1/6 ( n -1)

OpenStudy (trojanpoem):

Can you do it ? It's hard to do maths on screen.

OpenStudy (anonymous):

Waaaah.. ;~: I'm trying

OpenStudy (anonymous):

I want to say it's A ... But I'm probably wrong

OpenStudy (trojanpoem):

The arithmetic sequence won't work we don't add the same number each time.

OpenStudy (trojanpoem):

By the way, why don't you use the "dumbs way" ? Put 1 , 2, 3 ,4 the one which gives you the same numbers as the choices is the right one.

OpenStudy (trojanpoem):

B

OpenStudy (anonymous):

It's funny how the "dumbs way" helped me hahaha... Thank you for you assistance, sir SmartyPants. :)) <3

OpenStudy (trojanpoem):

Anytime.

OpenStudy (anonymous):

sorry, I couldn't remember the formulas for series/sequences lol

OpenStudy (trojanpoem):

Me too I think it can be solved using 

OpenStudy (anonymous):

\(\huge \Sigma\)